(25 points) Find the discrete time unit step response of the system with transfer function G(2) = 2(2+1)
Answer #1

Transfer function - Grz- ZCZ+1) Ycz)= 3-1 R(2) Z-1 Z L RC z czt)) ZCZ+0 YE RcZ) - Z Z-1 Yo-Z- Z ZCZ) EU Y(2= 7 Y(21= 2.8 (2+1